Can someone please help me with this equation?
The museum in Boston estimated that 64% of all visitors came from within the state. On Saturday 2,500 people attended the museum. How many attended the museum from out of state?
Thanks in advance for your help!

Answer by stanbon(75887)

x/2500 = 64/100
x = 0.64*2500
x = 1600
===============
Cheers,
Stan H.